What value of K will the roots of the equation 2x^2-3x+K=0 be imaginary?
Answer by Theo(13342)

equation is 2x^2 - 3x + k = 0

standard form of equation is ax^2 + bx + c = 0

in this equation:

a = 2
b = -3
c = k

discriminant is b^2 - 4ac

if the discriminant is negative, then the roots will be imaginary.

b^2 - 4ac becomes:

(-3)^2 - 4*2*k

this equals 9 - 8*k

solve for k to get:

k = 9/8

if k is greater than 9/8, then the discriminant will be negative and the roots will be imaginary.

example:

let k = 9/8

equation becomes:

2x^2 - 3x + 9/8 = 0

solve for the roots to get:

x = (-b +/- sqrt(b^2-4ac))/2a

a = 2
b = -3
c = 9/8

formula becomes:

x = -(9/8) +/- sqrt(9 - 4*2*9/8) / 4

this becomes:

x = (-9/8) +/- sqrt(9-9) / 4 = real roots.

now let c = 10/8

equation becomes:

x = (-10/8) +/- sqrt (9 - 4*2*10/8) / 4

this becomes:

x = (-10/8) +/- sqrt (9 - 10) / 4

sqrt (9-10) is negative causing the roots to be imaginary.

answer is that k > 9/8 makes the roots imaginary.